OpenStudy (anonymous):

How do you do this math problem? 14-(-6)=

OpenStudy (anonymous):

If you subtract a negative it's the same as adding so you would look at 14 + 6 = 20

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

14+6=20

OpenStudy (anonymous):

14-(-6)=14+6=20